OpenStudy (loganh):

Concrete can be purchased by the cubic yard. How much will it cost to pour a slab 18 feet by 18 feet by 4 inches for a patio if the concrete costs $41.00 per cubic yard?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (loganh):

$1476 is what I got

OpenStudy (skullpatrol):

Did you change all the units to yards?

OpenStudy (loganh):

no XD

OpenStudy (skullpatrol):



OpenStudy (mathmale):



OpenStudy (skullpatrol):

Workout the volume of the slab in yd^3

OpenStudy (mathmale):

Or find the volume in cubic feet and divide that result by \[\frac{ 27~ft^3 }{ 1~yd^3 }\]

OpenStudy (loganh):

164?

OpenStudy (loganh):

I divided 108 by 27

OpenStudy (mathmale):

Please show all steps of the work you've done, and please be certain to include the appropriate units of measurement.

OpenStudy (mathmale):

18 feet by 18 feet by 4 inches becomes (18 ft)(18 ft) ( [1/3] ft ) = ?

OpenStudy (loganh):

I typed that wrong sorry

OpenStudy (mathmale):

All I see that's wrong with your latest result is that you haven't included units of measurement.

OpenStudy (loganh):

$164

OpenStudy (loganh):

108/27 multiplied by 41

OpenStudy (loganh):

41 being $41

OpenStudy (mathmale):

Again I ask you to include units of measurement. $41 is incorrect; the correct version would be \[\frac{ $41 }{ yd^3 }\]

OpenStudy (loganh):

well i took the 108' and divided it by 27' to get the cube yards. I then multiplied this by $41- the price of a cubic yard of concrete

OpenStudy (mathmale):

toss out that $164. Multiply 108 yd^3 by \[\frac{ $41 }{ yd^3 }\] and simplify your result. What does this result represent?

OpenStudy (loganh):

its not 108 yd^3 it is 108'^3

OpenStudy (loganh):

I have to go. Thanks for the help

OpenStudy (mathmale):

YW